OpenStudy (anonymous):

the directions say "Find the ratio of girls to boys in a class, given the number of boys and the total number of students." ----What is 19 boys to 23 students?

OpenStudy (anonymous):

4:19

OpenStudy (hoblos):

number of girls = 23-19 =4 ratio = 4:19

OpenStudy (anonymous):

i accidently put 19 to 23. what i meant to put is 19 to 35 What is the answer?

OpenStudy (hoblos):

0k.. so number of girls =35-19= 16 ratio= 16:19

OpenStudy (anonymous):

thank you

OpenStudy (hoblos):

my pleasure XD

OpenStudy (anonymous):